# FedFACT: A Provable Framework for Controllable Group-Fairness Calibration in Federated Learning

### v1

@@ Abstract
With emerging application of Federated Learning (FL) in decision-making scenarios, it is imperative to regulate model fairness to prevent disparities across sensitive groups (e.g., female, male). Current research predominantly focuses on two concepts of group fairness within FL: Global Fairness (overall model disparity across all clients) and Local Fairness (the disparity within each client). However, the non-decomposable, non-differentiable nature of fairness criteria pose two fundamental, unresolved challenges for fair FL: (i) Harmonizing global and local fairness in multi-class classification; (ii) Enabling a controllable, optimal accuracy-fairness trade-off. To tackle the aforementioned challenges, we propose a novel controllable federated group-fairness calibration framework, named FedFACT. FedFACT identifies the Bayes-optimal classifiers under both global and local fairness constraints in multi-class case, yielding models with minimal performance decline while guaranteeing fairness. To effectively realize an adjustable, optimal accuracy-fairness balance, we derive specific characterizations of the Bayes-optimal fair classifiers for reformulating fair FL as personalized cost-sensitive learning problem for in-processing, and bi-level optimization for post-processing. Theoretically, we provide convergence and generalization guarantees for FedFACT to approach the near-optimal accuracy under given fairness levels. Extensive experiments on multiple datasets across various data heterogeneity demonstrate that FedFACT consistently outperforms baselines in balancing accuracy and global-local fairness.
@@ 1 Introduction
Federated learning (FL) is a collaborative distributed machine learning paradigm that allows multiple clients to jointly train a shared model while preserving the privacy of their local data fedavg . As FL is increasingly adopted in high-stakes domains—healthcare healthcare-1 ; healthcare-2 ; healthcare-3-nguyen2022federated , finance finance-PE-chouldechova2017fair ; finance-2-byrd2020differentially , and recommender systems recommendation-burke2017multisided ; recommendation-2-yang2020federated ; recommendation-3-sun2024survey —ensuring fairness is imperative to preventing discrimination against demographic groups based on sensitive attributes such as race, gender, age, etc. Although a rich literature addresses group fairness in centralized settings reduction-agarwal2018reductions ; beyond-alghamdi2022beyond ; pre-processing-jovanovic2023fare ; posthoc , these methods depend on full access to the entire dataset and centralized processing, imposing excessive communication overhead and elevating privacy concerns when directly applied in the FL context.
@@
However, existing methods face certain challenges in controlling group fairness in the federated environment: (i) Harmonizing global and local fairness in multi-class classification. Divergent sensitive-group distributions from client heterogeneity separate global and local fairness, thereby imposing an intrinsic trade-off AGLFOP-hamman2024demystifying ; cost-local-global-fair-duan2025cost . Most fair FL approaches focus exclusively on either global or local fairness fairfed ; FCFL-cui2021addressing ; fedfb-zeng2021improving ; agnostic-FL-du2021fairness ; Preference-Aware-federated-group-fairness , thereby inevitably sacrificing the other objective and impeding the realization of both fairness criteria. Moreover, these research predominantly addresses fairness in the binary case, despite the ubiquity of multiclass tasks in practical FL scenarios. (ii) Enabling a controllable, optimal accuracy-fairness trade-off with theoretical guarantee. The non-decomposable, non-differentiable nature of group-fairness measures poses significant optimization challenges fair-survey-mehrabi2021survey ; jmlr-recall-cotter2019optimization . Existing federated frameworks typically rely on surrogate fairness losses Preference-Aware-federated-group-fairness ; fairfed ; fedfb-zeng2021improving ; FCFL-cui2021addressing ; global-local-EquiFL-makhija2024achieving , yet the inevitable surrogate-fairness gap surrogate-fairness-gap-yao2024understanding ; too-relax-pmlr-v119-lohaus20a produces suboptimal performance and undermines convergence stability, thus hampering the controllability of the accuracy-fairness trade-off.
@@
To address these challenges, we propose a novel Federated group-FAirness CalibraTion framework, named FedFACT, comprising in-processing and post-processing approaches. Our framework is capable of ensuring controllable global and local fairness with minimal accuracy deterioration, underpinned by provable convergence and consistency guarantees. To harmonize global and local fairness, we seek to find the optimal classifier under both dual fairness constraints in the multi-class case. To this end, specific characterizations of the federated Bayes-optimal fair classifiers are established for both the in-processing and post-processing phases in FL. Building on the Bayes-optimal fair classifier’s structure, we develop efficient, privacy-preserving federated optimization strategies that realize a controllable and theoretically optimal fairness–accuracy trade-off. In detail, FedFACT reduces the in-processing task into a series of personalized cost-sensitive classification problems, and reformulates post-processing as a bi-level optimization that leverages the explicit form of the federated Bayes-optimal fair classifiers. We further derive theoretical convergence and generalization guarantees, demonstrating that our methods achieve near-optimal model performance while enforcing tunable global and local fairness constraints.
@@
Our extensive experiments on multiple real-world datasets verify the efficiency and effectiveness of FedFACT, highlighting that FedFACT delivers a superior, controllable accuracy-fairness balance while maintaining competitive classification performance compared to state-of-the-art methods. Our main contributions are summarized as follows:
@@ 2 Related Work
Group Fairness in Machine Learning. As summarized in previous work fair-survey-mehrabi2021survey , group fairness is broadly defined as the absence of prejudice or favoritism toward a sensitive group based on their inherent characteristics. Common strategies for realizing group fairness in machine learning can be classified into three categories: pre-, in-, and post-processing methods. Pre-processing fair-pre-pmlr-v162-li22p ; fair-pre-2-pmlr-v202-jovanovic23a ; fair-pre-3-10.1145/3394486.3403080 approaches aim to modify training data to eradicate underlying bias before model training. In-processing inpro-pmlr-v162-kim22b ; inpro-pmlr-v202-li23h ; reduction-agarwal2018reductions ; overlap ; surrogate-fairness-gap-yao2024understanding ; inpro-fairness-nips2024-1-zeng2024fairness ; inpro-fairness-nips2024-2-yazdani2024fair methods are developed to achieve fairness requirements by intervention during the training process. Post-processing posthoc ; fair-guarantees-demographic-denis2024fairness ; bayes-2-zeng2024bayes ; fair-and-optimal-postpro-pmlr-v202-xian23b ; unified-postprocessing-framework-group-xian2024 ; EO-hardt2016equality ; FACT-confusion methods adjust the prediction results generated by a given model to adapt to fairness constraints after training stage. However, because these methods require access to the full dataset, they are confined to mitigating disparities only at the local level.
@@
Furthermore, previous work AGLFOP-hamman2024demystifying offered a theoretical study elucidating the divergence between local and global fairness in FL, while revealing the intrinsic trade-off between these fairness objectives and accuracy. cost-local-global-fair-duan2025cost formulates local and global fairness optimization into a linear program for minimal fairness cost, but it does not realize the Bayes-optimal balance between accuracy and fairness. logofair-zhang2025logofair  derives the Bayes-optimal classifier and decomposes the overall problem into client-specific optimizations, yet their approach applies only to binary classification in post-processing. Persistent challenges include inadequate accuracy-fairness flexibility and limited convergence guarantee in mitigating disparities within multi-class classification across various stages of FL training.
@@ 3 Preliminary
Notation. Denote by $\mathbf{1}_{m}$ the $m$-dimensional all-ones vector and by $\mathbf{1}_{m\times m}$ the $m\times m$ all-ones matrix. Write the probability simplex as $\Delta_{m}=\{\,q\in[0,1]^{m}:\|q\|_{1}=1\}$, and let $e_{i}$ be the $i$-th standard basis vector. Throughout this paper, bold lowercase letters denote vectors and bold uppercase letters denote matrices. For two equally sized matrices $\mathbf{A}$ and $\mathbf{B}$, their Frobenius inner product is $\langle\mathbf{A},\mathbf{B}\rangle=\sum_{i,j}a_{ij}\,b_{ij}$. For positive integer $n$, $[n]=\{1,\dots,n\}$.
Fairness in classification. Let $(X,A,Y)$ be a random tuple, where $X\in\mathcal{X}$ for some feature space $\mathcal{X}\subseteq\mathbb{R}^{d}$, labels $Y\in\mathcal{Y}=[m]$, and the discrete sensitive attribute $A\in\mathcal{A}$. Given the randomized classifiers ${h}:\mathcal{X}\times\mathcal{A}\to\Delta_{m}$, the prediction $\widehat{Y}$ is associated with the random outputs of ${h}$ defined by $\mathbb{P}(\widehat{Y}=y\mid\mathbf{x})={h}_{y}(\mathbf{x})$. In this work, we generally focus on three popular group-fairness criteria—Demographic Parity (DP) DP-dwork2012fairness , Equalized Opportunity (EOP) EO-hardt2016equality and Equalized Odds EO-hardt2016equality —in multiclass classification tasks with multiple sensitive attributes, as defined in prior works fair-guarantees-demographic-denis2024fairness ; fair-and-optimal-postpro-pmlr-v202-xian23b ; unified-postprocessing-framework-group-xian2024 .
@@
Matrix Notations ($\begin{array}[]{l}\text{Matrix Notations }\\ \text{(Centralized)}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{i}\big{(}\mathbb{I}[a=a^{\prime}]-p_{a}\big{)}\mathbf{C}_{i,y}^{a}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\big{(}\frac{p_{a^{\prime}}}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{1}{p_{y}}\big{)}\mathbf{C}_{y,y}^{a}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$
@@
Global Fairness ($\begin{array}[]{l}\text{Global Fairness }\\ \text{(Federated)}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{k}\sum_{i}\big{(}p_{k|a^{\prime}}\mathbb{I}[a=a^{\prime}]-p_{a,k}\big{)}\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{k}\big{(}\frac{p_{a^{\prime},k}}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y}}\big{)}\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$
@@
Local Fairness ($\begin{array}[]{l}\text{Local Fairness }\\ \text{(Federated)}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{i}\big{(}\mathbb{I}[a=a^{\prime}]-p_{a|k}\big{)}\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{k}\big{(}\frac{p_{a^{\prime},k}}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}\big{)}\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$
@@
Fairness and performance metrics. As presented in Table 1 (EO criterion and notation explanations see Appendix A), the global fairness constraints typically can be expressed by $|\mathscr{D}^{g}_{u_{g}}(h)|\leq\xi^{g},u_{g}\in\mathcal{U}_{g}$, where $\mathscr{D}^{g}_{u_{g}}(h)=\sum_{a}\sum_{k}\langle\mathbf{D}^{a,k}_{u_{g}},\mathbf{C}^{a,k}(h)\rangle$ represents the constraints required to achieve the global fairness criterion. Similarly, the local fairness constraints are denoted as $|\mathscr{D}^{l,k}_{u_{l,k}}(h)|\leq\xi^{l,k},u_{l,k}\in\mathcal{U}_{l,k}$, with $\mathscr{D}^{l,k}_{u_{l,k}}(h)=\sum_{a}\langle\mathbf{D}^{a,k}_{u_{l,k}},\mathbf{C}^{a,k}(h)\rangle$. For performance metrics, we consider a risk metric expressed as a linear function of the population confusion matrix, i.e. $\mathcal{R}(h)=\langle\mathbf{R},\mathbf{C}(h)\rangle=\sum_{a}\sum_{k}p_{a,k}\langle\mathbf{R},\mathbf{C}^{a,k}(h)\rangle$. This formulation has been explored in multi-label and fair classification contexts consistent-multilabel-classification-wang2019consistentclassificationgeneralizedmetrics ; overlap , and encompasses a variety of performance metrics, such as average recall and precision. In this paper, we primarily focus on standard classification error to set $\mathbf{R}=\mathbf{1}_{m\times m}-\mathbf{I}$.

### v2

@@ Abstract
With the emerging application of Federated Learning (FL) in decision-making scenarios, it is imperative to regulate model fairness to prevent disparities across sensitive groups (e.g., female, male). Current research predominantly focuses on two concepts of group fairness within FL: Global Fairness (overall model disparity across all clients) and Local Fairness (the disparity within each client). However, the non-decomposable, non-differentiable nature of fairness criteria poses two fundamental, unresolved challenges for fair FL: (i) Harmonizing global and local fairness, especially in multi-class setting; (ii) Enabling a controllable, optimal accuracy-fairness trade-off. To tackle these challenges, we propose a novel controllable federated group-fairness calibration framework, named FedFACT. FedFACT identifies the Bayes-optimal classifiers under both global and local fairness constraints, yielding models with minimal performance decline while guaranteeing fairness. Building on the characterization of the optimal fair classifiers, we reformulate fair federated learning as a personalized cost-sensitive learning problem for in-processing and a bi-level optimization for post-processing. Theoretically, we provide convergence and generalization guarantees for FedFACT to approach the near-optimal accuracy under given fairness levels. Extensive experiments on multiple datasets across various data heterogeneity demonstrate that FedFACT consistently outperforms baselines in balancing accuracy and global-local fairness.
@@ 1 Introduction
Federated learning (FL) is a collaborative distributed machine learning paradigm that allows multiple clients to jointly train a shared model while preserving the privacy of their local data fedavg. As FL is increasingly adopted in high-stakes domains—healthcare healthcare-1; healthcare-2; healthcare-3-nguyen2022federated, finance finance-PE-chouldechova2017fair; finance-2-byrd2020differentially; ref-1, pattern recognition pat-1; pat-2; ref-2; pattern-recognition-3; pat-3, and recommender systems recommendation-burke2017multisided; recommendation-2-yang2020federated; han2024hypergraph; recommendation-3-sun2024survey—ensuring fairness is imperative to prevent discrimination against demographic groups based on sensitive attributes pattern-recognition-1; pattern-recognition-2; ForgetMe; Speech-bias, such as race, gender, age, etc. Although a rich literature addresses group fairness in centralized settings reduction-agarwal2018reductions; beyond-alghamdi2022beyond; pre-processing-jovanovic2023fare; posthoc, these methods depend on full access to the entire dataset and centralized processing, imposing excessive communication overhead and elevating privacy concerns when directly applied in the FL context.
@@
However, existing methods face certain challenges in controlling group fairness within FL: (i) Harmonizing global and local fairness, especially in multi-class classification. Divergent sensitive-group distributions from client heterogeneity separate global and local fairness, thereby imposing an intrinsic trade-off AGLFOP-hamman2024demystifying; cost-local-global-fair-duan2025cost. Most fair FL approaches focus exclusively on either global or local fairness fairfed; FCFL-cui2021addressing; fedfb-zeng2021improving; agnostic-FL-du2021fairness; Preference-Aware-federated-group-fairness, thereby inevitably sacrificing the other objective and impeding the realization of both fairness criteria. Moreover, this research predominantly addresses fairness in the binary case, despite the ubiquity of multiclass tasks in practical FL scenarios. (ii) Enabling a controllable, optimal accuracy-fairness trade-off with theoretical guarantee. The non-decomposable, non-differentiable nature of group-fairness measures poses significant optimization challenges fair-survey-mehrabi2021survey; jmlr-recall-cotter2019optimization. Existing frameworks typically rely on surrogate fairness losses Preference-Aware-federated-group-fairness; fairfed; fedfb-zeng2021improving; FCFL-cui2021addressing; global-local-EquiFL-makhija2024achieving, yet the inevitable surrogate-fairness gap surrogate-fairness-gap-yao2024understanding; too-relax-pmlr-v119-lohaus20a produces suboptimal performance and undermines convergence stability, thus hampering the controllability of the accuracy-fairness trade-off.
@@
To address these challenges, we propose a novel Federated group-FAirness CalibraTion framework, named FedFACT, comprising in-processing and post-processing approaches. Our framework is capable of ensuring controllable global and local fairness with minimal accuracy deterioration, underpinned by provable convergence and consistency guarantees. To harmonize global and local fairness, we seek to find the optimal classifier under both dual fairness constraints in the multi-class case. To this end, specific characterizations of the federated Bayes-optimal fair classifiers are established for both the in‐processing and post‐processing phases in FL. Building on the Bayes-optimal fair classifier’s structure, we develop efficient, privacy-preserving federated optimization strategies that realize a controllable and theoretically optimal fairness–accuracy trade-off. In detail, FedFACT reduces the in-processing task into a series of personalized cost-sensitive classification problems, and reformulates post-processing as a bi-level optimization that leverages the explicit form of the federated Bayes-optimal fair classifiers. We further derive theoretical convergence and generalization guarantees, demonstrating that our methods achieve near-optimal model performance while enforcing tunable global and local fairness constraints.
@@
Our extensive experiments on multiple real-world datasets verify the efficiency and effectiveness of FedFACT, highlighting that FedFACT delivers a superior, controllable accuracy-fairness balance while maintaining competitive classification performance compared to state-of-the-art methods.
Our main contributions are summarized as follows:
@@ 2 Related Work
Group Fairness in Machine Learning. As summarized in previous work fair-survey-mehrabi2021survey, group fairness is broadly defined as the absence of prejudice or favoritism toward a sensitive group based on their inherent characteristics. Common strategies for realizing group fairness in machine learning can be classified into three categories: pre-, in-, and post-processing methods. Pre-processing fair-pre-pmlr-v162-li22p; fair-pre-2-pmlr-v202-jovanovic23a; fair-pre-3-10.1145/3394486.3403080 approaches aim to modify training data to eradicate underlying bias before model training. In-processing inpro-pmlr-v162-kim22b; inpro-pmlr-v202-li23h; reduction-agarwal2018reductions; overlap; surrogate-fairness-gap-yao2024understanding; Physics-constrained-symbolic-regression; inpro-fairness-nips2024-1-zeng2024fairness; inpro-fairness-nips2024-2-yazdani2024fair methods are developed to achieve fairness requirements by intervening during the training process. Post-processing posthoc; fair-guarantees-demographic-denis2024fairness; bayes-2-zeng2024bayes; fair-and-optimal-postpro-pmlr-v202-xian23b; unified-postprocessing-framework-group-xian2024; EO-hardt2016equality; FACT-confusion methods adjust the prediction results generated by a given model to adapt to fairness constraints after the training stage. However, because these methods require access to the full dataset, they are confined to mitigating disparities only at the local level.
@@
Furthermore, previous work AGLFOP-hamman2024demystifying offered a theoretical study elucidating the divergence between local and global fairness in FL, while revealing the intrinsic trade-off between these fairness objectives and accuracy. cost-local-global-fair-duan2025cost formulates local and global fairness optimization into a linear program for minimal fairness cost, but it does not realize the Bayes-optimal balance between accuracy and fairness. logofair-zhang2025logofair derives the Bayes-optimal classifier and decomposes the overall problem into client-specific optimizations, yet their approach applies only to binary classification in post-processing. Persistent challenges include inadequate accuracy-fairness flexibility and limited convergence guarantee in mitigating disparities within multi‐class classification across various stages of FL training.
@@ 3 Preliminary
Notation. Denote by $\mathbf{1}_{m}$ the $m$-dimensional all-ones vector and by $\mathbf{1}_{m\times m}$ the $m\times m$ all-ones matrix. Write the probability simplex as $\Delta_{m}=\{\,q\in[0,1]^{m}:\|q\|_{1}=1\}$, and let $e_{i}$ be the $i$-th standard basis vector. Throughout this paper, bold lowercase letters denote vectors and bold uppercase letters denote matrices. For two equally sized matrices $\mathbf{A}$ and $\mathbf{B}$, their Frobenius inner product is $\langle\mathbf{A},\mathbf{B}\rangle=\sum_{i,j}a_{ij}\,b_{ij}$. For positive integer $n$, $n]=\{1,\dots,n\}$.
Fairness in classification. Let $(X,A,Y)$ be a random tuple, where $X\in\mathcal{X}$ for some feature space $\mathcal{X}\subseteq\mathbb{R}^{d}$, labels $Y\in\mathcal{Y}=[m]$, and the discrete sensitive attribute $A\in\mathcal{A}$. Given the randomized classifiers ${h}:\mathcal{X}\times\mathcal{A}\to\Delta_{m}$, the prediction $\widehat{Y}$ is associated with the random outputs of h{h} defined by $\mathbb{P}(\widehat{Y}=y\mid\mathbf{x})={h}_{y}(\mathbf{x})$. In this work, we generally focus on three popular group-fairness criteria—Demographic Parity (DP) DP-dwork2012fairness, Equalized Opportunity (EOP) EO-hardt2016equality and Equalized Odds EO-hardt2016equality—in multiclass classification tasks with multiple sensitive attributes, as defined in prior works fair-guarantees-demographic-denis2024fairness; fair-and-optimal-postpro-pmlr-v202-xian23b; unified-postprocessing-framework-group-xian2024.
@@
$\begin{array}[]{l}\text{ Group Constraints }\\ \text{(Centralized)}\end{array}$ $\begin{array}[]{l}\left|\mathbb{P}(\widehat{Y}=y\mid A=a^{\prime})-\mathbb{P}(\widehat{Y}=y)\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$ $\begin{array}[]{l}\left|\mathbb{P}(\widehat{Y}=y\mid A=a^{\prime},Y=y)-\mathbb{P}(\widehat{Y}=y\mid Y=y)\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$
@@
$\begin{array}[]{l}\text{Matrix Notations }\\ \text{(Centralized)}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{i}\big(\mathbb{I}[a=a^{\prime}]-p_{a}\big)\mathbf{C}_{i,y}^{a}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\big(\frac{p_{a^{\prime}}}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{1}{p_{y}}\big)\mathbf{C}_{y,y}^{a}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$
@@
$\begin{array}[]{l}\text{Global Fairness }\\ \text{(Federated)}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{k}\sum_{i}\big(p_{k|a^{\prime}}\mathbb{I}[a=a^{\prime}]-p_{a,k}\big)\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{k}\big(\frac{p_{a^{\prime},k}}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y}}\big)\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$
$\begin{array}[]{l}\text{Local Fairness }\\ \text{(Federated)}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{i}\big(\mathbb{I}[a=a^{\prime}]-p_{a|k}\big)\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$ $\begin{array}[]{l}\left|\sum_{a}\sum_{k}\big(\frac{p_{a^{\prime},k}}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}\big)\mathbf{C}^{a,k}_{i,y}\right|\\ \forall a^{\prime}\in\mathcal{A},\forall y\in\mathcal{Y}\end{array}$
@@
Fairness and performance metrics. As presented in Table 1 (EO criterion and notation explanations see Appendix A), the global fairness constraints typically can be expressed by $|\mathscr{D}^{g}_{u_{g}}(h)|\leq\xi^{g},u_{g}\in\mathcal{U}_{g}$, where $\mathscr{D}^{g}_{u_{g}}(h)=\sum_{a}\sum_{k}\langle\mathbf{D}^{a,k}_{u_{g}},\mathbf{C}^{a,k}(h)\rangle$ represents the constraints required to achieve the global fairness criterion. Similarly, the local fairness constraints are $|\mathscr{D}^{k}_{u_{k}}(h)|\leq\xi^{k},u_{k}\in\mathcal{U}_{k},k\in[N]$, with $\mathscr{D}^{k}_{u_{k}}(h)=\sum_{a}\langle\mathbf{D}^{a,k}_{u_{k}},\mathbf{C}^{a,k}(h)\rangle$. For performance metrics, we consider a risk metric expressed as a linear function of the population confusion matrix, i.e. $\mathcal{R}(h)=\langle\mathbf{R},\mathbf{C}(h)\rangle=\sum_{a}\sum_{k}p_{a,k}\langle\mathbf{R},\mathbf{C}^{a,k}(h)\rangle$. This formulation has been explored in multi-label and fair classification contexts consistent-multilabel-classification-wang2019consistentclassificationgeneralizedmetrics; overlap, and encompasses a variety of performance metrics, such as average recall and precision. In this paper, we primarily focus on standard classification error to set $\mathbf{R}=\mathbf{1}_{m\times m}-\mathbf{I}$.